# Multi-Modal Data-Efficient 3D Scene Understanding for Autonomous Driving

### v2

@@ 4 LaserMix++
To leverage interactions between the camera and LiDAR, we first elaborate on the correspondence between LiDAR and cameras (Sec. 4.1). We then propose a multi-modal LaserMix operation (Sec. 4.2), a cross-sensor feature distillation (Sec. 4.3), and language-driven knowledge guidance (Sec. 4.4). Finally, we describe the overall LaserMix++ pipeline for enhancing data-efficient 3D scene understanding (Sec. 4.5).
Prevailing driving perception systems often consist of a spectrum of sensors to ensure safe operations [26, 27, 28]. A typical sensor setup incorporated by existing perception systems contains both LiDAR and surrounding cameras. Both 2D (trained w/ RGB images) and 3D (trained w/ LiDAR point clouds) perception models are leveraged which can complement each other, especially in challenging conditions such as low light, adverse weather, and motion perturbations [29, 30, 31, 32]. In this work, instead of solely relying on LiDAR point clouds to train semi-supervised learning models, we propose to leverage the abundant RGB images from cameras as additional resources for a more holistic multi-modal data-efficient 3D scene understanding. This enhanced framework, dubbed LaserMix++, does not require any image annotations, thus maintaining the same data efficiency as the baseline framework described in Sec. 3.3.
@@
$[p^{u},p^{v},1]^{\text{T}}=\frac{1}{p^{z}}\times\Gamma_{K}\times\Gamma_{c\leftarrow l}\times[p^{x},p^{y},p^{z}]^{\text{T}}\,,$ (10)
@@
It is worth mentioning that the correspondences between points and pixels do not always hold due to the possible mis-overlap between the field-of-views of the LiDAR and the camera. To handle this, we generate a correspondence mask $\mathcal{M}$ to establish valid point-pixel relationships. The process begins by projecting LiDAR points onto the image plane using the known extrinsic and intrinsic camera parameters, allowing us to determine which LiDAR points have corresponding pixels in the image. If a LiDAR point $x_{\text{p}}$ projects onto a valid pixel location, the corresponding entry in the mask $\mathcal{M}$ is set to $1$. For LiDAR points that project outside the image boundaries or onto invalid regions, the corresponding entry in $\mathcal{M}$ is set to 00. These entries are then padded with zeros in the pixel correspondence data. This automated process leverages existing geometric transformations and does not require additional manual data annotation, ensuring scalability and efficiency.
@@ 4.2 Multi-Modal LaserMix
Different from the pure positional information encoded in the LiDAR point cloud, the image pixels provide extra texture information that could be supplementary to the 3D scene understanding task. To leverage such visual guidance for data-efficient learning, we associate the image pixels with the LiDAR points during LaserMix. Similar to Equation 7, such a multi-modal data mixing process between the LiDAR and camera can be defined as:
$\begin{split}\tilde{\sigma}_{1},\tilde{\sigma}_{2}&=\textrm{Multi-Modal LaserMix}[\{x_{\text{p}},x_{\text{img}}\}_{1},\{x_{\text{p}},x_{\text{img}}\}_{2}]\,,\\ \tilde{\sigma}_{1}&=\{x_{\text{p}},x_{\text{img}}\}_{1}^{a_{1}}\cup\{x_{\text{p}},x_{\text{img}}\}_{2}^{a_{2}}\cup\{x_{\text{p}},x_{\text{img}}\}_{1}^{a_{3}}\cup\cdots,\\ \tilde{\sigma}_{2}&=\{x_{\text{p}},x_{\text{img}}\}_{2}^{a_{1}}\cup\{x_{\text{p}},x_{\text{img}}\}_{1}^{a_{2}}\cup\{x_{\text{p}},x_{\text{img}}\}_{2}^{a_{3}}\cup\cdots.\\ \end{split}$ (11)
The area set $\{a_{1},a_{2},...,a_{m}\}$ can be obtained in the same way as Sec. 3.2. The multi-modal variant of LaserMix leverages off-the-shelf visual information from synchronized camera images to assist the consistency regularization in Sec. 3.3. Such an approach further enhances the effect of spatial priors in driving scenes, since the LiDAR points have been “painted” with extra texture information from the images. The use of the correspondence mask $\mathcal{M}$ ensures that only valid LiDAR point-pixel pairs are utilized, maintaining the consistency of the multi-modal information. After obtaining mixed LiDAR-image pairs, we propose the following two operations in the LaserMix++ framework to assist the data-efficient 3D scene understanding task with such multi-modal inputs.
@@ 4.3 Camera-to-LiDAR Feature Distillation
$\begin{split}F_{\text{p}}=\mathcal{H}_{\zeta}^{s}(\hat{\mathcal{G}}_{\hat{\theta}}^{s}(x_{\text{p}}))\,,~{}~{}~{}F_{\text{img}}=\hat{\mathcal{G}}_{\xi}^{\text{img}}(x_{\text{img}})\,,\end{split}$ (12)
where $\hat{\mathcal{G}}_{\hat{\theta}}$ with parameters $\hat{\theta}$ is the backbone of the Student network $\mathcal{G}_{\theta}^{t}$, and $\mathcal{H}_{\zeta}^{s}$ is a projection layer mapping the dimension of point cloud features to match that of image features. Based on the sensor calibration parameters from Equation 10, we align the features to create paired point-image features $\{F_{\text{p}},\tilde{F}_{\text{img}}\}$. A key aspect of our approach is ensuring that the integration of image-based semantics does not diminish the unique spatial features of the point cloud.
@@
where $\langle$,$\rangle$ , ⟩ calculates the inner product. The use of cosine distance ensures that the semantic alignment process enriches the point cloud features without overwhelming them, thereby achieving a balanced integration of both semantic and geometric information. This approach facilitates the transfer of semantically coherent features from images to LiDAR, enhancing the LiDAR segmentation model’s performance in a data-efficient manner.
@@ 4.4 Language-Driven Knowledge Guidance
Given class names as text prompts, we use the CLIP text encoder $\mathcal{G}^{\text{txt}}_{\varrho}$ to extract text embeddings $F^{{}^{\prime}}_{\text{txt}}$, parameterized by $\varrho$. These embeddings, combined with the image features $F_{\text{img}}$ from Equation 12, are processed through the pre-trained open-vocabulary image segmentation head $\mathcal{H}_{\varsigma}^{\text{img}}$ and the LiDAR segmentation model $\mathcal{G}_{\theta}^{s}$:
$\begin{split}F^{{}^{\prime}}_{\text{p}}=\mathcal{G}_{\theta}^{s}(x_{\text{p}})\,,~{}~{}~{}F^{{\prime}}_{\text{img}}=\mathcal{H}{\varsigma}^{\text{img}}(F_{\text{img}})\circ F^{{\prime}}_{\text{txt}}\,,\end{split}$ (14)
@@
Role in our framework: Our enhanced framework is designed to leverage both LiDAR and image data for data-efficient 3D scene understanding, without needing image labels from the target driving datasets. We first propose multi-modal LaserMix to enable richer interactions across sensors. To align point cloud and image features, we adapt a pretrained image segmentation model to the LiDAR segmentation backbone. To obtain auxiliary supervision signals for unlabeled scans, we propose to generate text-aligned non-probabilistic outputs from images and match them with that of the LiDAR point clouds. As we will show in the following sections, these two modules contribute to significant performance improvements for data-efficient 3D scene understanding.
@@ 4.5 Overall Framework
Incorporating everything together, we now present the overall LaserMix++ framework. Based on the 3D scene consistency regularization baseline in Sec. 3.3, our approach seamlessly integrates optimization of objectives as follows:
  • The conventional supervision signals from labeled data, as in Equation 8.
  • The mixing-based cross-sensor consistency from multi-modal LaserMix, as in Equation 11.
  • The consistency regularization between Student and Teacher networks, as in Equation 9.
  • The camera-to-LiDAR feature distillation between images and LiDAR point clouds, as in Equation 13.
  • The auxiliary supervisions for unlabeled data, obtained from text-driven knowledge guidance in Equation 15.
@@ 5 Experiments
[…]
@@
Range View Sup.-only - FIDNet $38.3$ $57.5$ $62.7$ $67.6$ $36.2$ $52.2$ $55.9$ $57.2$ $33.1$ $47.7$ $49.9$ $52.5$
@@
Voxel Sup.-only - Cylinder3D $50.9$ $65.9$ $66.6$ $71.2$ $45.4$ $56.1$ $57.8$ $58.7$ $39.2$ $48.0$ $52.1$ $53.8$
@@
LiM3D [11] CVPR’23 - - - - - $61.6$ $62.6$ $62.8$ $60.3$ $60.5$ $60.9$
ImageTo360 [99] ICCVW’23 - - - - $54.1$ $60.0$ $62.2$ $65.0$
@@
Fusion Sup.-only - SPVCNN $57.9$ $71.7$ $73.0$ $74.6$ $52.7$ $64.1$ $64.5$ $65.1$ $47.2$ $57.3$ $58.2$ $58.8$
@@
[…]
@@
SemanticKITTI-C None - MinkUNet $100.0$ $81.9$ $55.9$ $54.0$ $53.3$ $32.9$ $56.3$ $58.3$ $54.4$ $46.1$
@@
[…]
# Method $\mathbf{1/16}$ $\mathbf{1/8}$ $\mathbf{1/4}$ $\mathbf{1/2}$
a MeanTeacher [18] $66.1$ $71.2$ $74.4$ $76.3$
CCT [19] $66.4$ $72.5$ $75.7$ $76.8$
GCT [20] $65.8$ $71.3$ $75.3$ $77.1$
CPS [21] $69.8$ $74.4$ $76.9$ $78.6$
b CPS-CutMix [21] $74.5$ $76.6$ $77.8$ $78.8$
[…]
# $\mathcal{L}_{\text{mt}}$ $\mathcal{L}_{\text{mix}}$ SS TS $\mathcal{L}_{\text{c2l}}$ $\mathcal{L}_{\text{lkg}}$ SemanticKITTI [28] nuScenes [38]
@@
TABLE VII: Ablation study on laser beam partition strategies in LaserMix. Horizontal axis: inclination direction $\phi$; Vertical axis: azimuth direction $\alpha$. A ($i$-$\alpha$, $j$-$\phi$) strategy denotes that there are $i$ azimuth and $j$ inclination partitions in total.
@@
$60.4$ $63.5_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+3.1)}}}}$ $65.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+4.8)}}}}$ $66.5_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+6.1)}}}}$ $66.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.8)}}}}$ $65.4_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.0)}}}}$
@@
$61.5_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+1.1)}}}}$ $63.3_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+2.9)}}}}$ $65.9_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.5)}}}}$ $66.1_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.7)}}}}$ $66.7_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+6.3)}}}}$ $65.3_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+4.9)}}}}$
@@
$60.9_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+0.6)}}}}$ $64.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+3.8)}}}}$ $65.9_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.5)}}}}$ $66.3_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.9)}}}}$ $66.0_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.6)}}}}$ $65.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+4.8)}}}}$
@@
[…]

### v3

@@ 4 LaserMix++
To leverage interactions between the camera and LiDAR, we first elaborate on the correspondence between LiDAR and cameras (Section 4.1). We then propose a multi-modal LaserMix operation (Section 4.2), a cross-sensor feature distillation (Section 4.3), and language-driven knowledge guidance (Section 4.4). Finally, we describe the overall LaserMix++ pipeline for enhancing data-efficient 3D scene understanding (Section 4.5).
Prevailing driving perception systems often consist of a spectrum of sensors to ensure safe operations [5, 80, 2]. A typical sensor setup incorporated by existing perception systems contains both LiDAR and surrounding cameras. Both 2D (trained w/ RGB images) and 3D (trained w/ LiDAR point clouds) perception models are leveraged, which can complement each other, especially in challenging conditions such as low light, adverse weather, and motion perturbations [94, 11, 36, 87]. In this work, instead of solely relying on LiDAR point clouds to train semi-supervised learning models, we propose to leverage the abundant RGB images from cameras as additional resources for a more holistic, multi-modal data-efficient 3D scene understanding. This enhanced framework, dubbed LaserMix++, does not require any image annotations, thus maintaining the same data efficiency as the baseline framework described in Section 3.3.
@@
$p^{u},p^{v},1]^{\text{T}}=\frac{1}{p^{z}}\times\Gamma_{K}\times\Gamma_{c\leftarrow l}\times[p^{x},p^{y},p^{z}]^{\text{T}}\,$, (10)
@@
It is worth mentioning that the correspondences between points and pixels do not always hold due to the possible mis-overlap between the field-of-view of the LiDAR and the camera. To handle this, we generate a correspondence mask $\mathcal{M}$ to establish valid point-pixel relationships. The process begins by projecting LiDAR points onto the image plane using the known extrinsic and intrinsic camera parameters, allowing us to determine which LiDAR points have corresponding pixels in the image. If a LiDAR point $x_{\text{p}}$ projects onto a valid pixel location, the corresponding entry in the mask $\mathcal{M}$ is set to $1$. For LiDAR points that project outside the image boundaries or onto invalid regions, the corresponding entry in $\mathcal{M}$ is set to 0. These entries are then padded with zeros in the pixel correspondence data. This automated process leverages existing geometric transformations and does not require additional manual data annotation, ensuring scalability and efficiency.
@@ 4.2 Multi-Modal LaserMix
Different from the pure positional information encoded in the LiDAR point cloud, the image pixels provide extra texture information that could be supplementary to the 3D scene understanding task. To leverage such visual guidance for data-efficient learning, we associate the image pixels with the LiDAR points during LaserMix. Similar to Eq. 7, such a multi-modal data mixing process between the LiDAR and camera can be defined as:
[…]
The area set $\{a_{1},a_{2},...,a_{m}\}$ can be obtained in the same way as Section 3.2. The multi-modal variant of LaserMix leverages off-the-shelf visual information from synchronized camera images to assist the consistency regularization in Section 3.3. Such an approach further enhances the effect of spatial priors in driving scenes, since the LiDAR points have been “painted” with extra texture information from the images. The use of the correspondence mask $\mathcal{M}$ ensures that only valid LiDAR point-pixel pairs are utilized, maintaining the consistency of the multi-modal information. After obtaining mixed LiDAR-image pairs, we propose the following two operations in the LaserMix++ framework to assist the data-efficient 3D scene understanding task with such multi-modal inputs.
@@ 4.3 Camera-to-LiDAR Feature Distillation
[…]
where $\hat{\mathcal{G}}_{\hat{\theta}}$ with parameters $\hat{\theta}$ is the backbone of the Student network $\mathcal{G}_{\theta}^{t}$, and $\mathcal{H}_{\zeta}^{s}$ is a projection layer mapping the dimension of point cloud features to match that of image features. Based on the sensor calibration parameters from Eq. 10, we align the features to create paired point-image features $\{F_{\text{p}},\tilde{F}_{\text{img}}\}$. A key aspect of our approach is ensuring that the integration of image-based semantics does not diminish the unique spatial features of the point cloud.
@@
where $\langle,\rangle$ calculates the inner product. The use of cosine distance ensures that the semantic alignment process enriches the point cloud features without overwhelming them, thereby achieving a balanced integration of both semantic and geometric information. This approach facilitates the transfer of semantically coherent features from images to LiDAR, enhancing the LiDAR segmentation model’s performance in a data-efficient manner.
@@ 4.4 Language-Driven Knowledge Guidance
Given class names as text prompts, we use the CLIP text encoder $\mathcal{G}^{\text{txt}}_{\varrho}$ to extract text embeddings $F^{{}^{\prime}}_{\text{txt}}$, parameterized by $\varrho$. These embeddings, combined with the image features $F_{\text{img}}$ from Eq. 12, are processed through the pre-trained open-vocabulary image segmentation head $\mathcal{H}_{\varsigma}^{\text{img}}$ and the LiDAR segmentation model $\mathcal{G}_{\theta}^{s}$:
[…]
@@
Role in our framework: Our enhanced framework is designed to leverage both LiDAR and image data for data-efficient 3D scene understanding, without needing image labels from the target driving datasets. We first propose the multi-modal LaserMix to enable richer interactions across sensors. To align point cloud and image features, we adapt a pretrained image segmentation model to the LiDAR segmentation backbone. To obtain auxiliary supervision signals for unlabeled scans, we propose to generate text-aligned non-probabilistic outputs from images and match them with those of the LiDAR point clouds. As we will show in the following sections, these two modules contribute to significant performance improvements for data-efficient 3D scene understanding.
@@ 4.5 Overall Framework
Incorporating everything together, we now present the overall LaserMix++ framework. Based on the 3D scene consistency regularization baseline in Section 3.3, our approach seamlessly integrates optimization of objectives as follows:
  • The conventional supervision signals from labeled data, as in Eq. 8.
  • The mixing-based cross-sensor consistency from multi-modal LaserMix, as in Eq. 11.
  • The consistency regularization between Student and Teacher networks, as in Eq. 9.
  • The camera-to-LiDAR feature distillation between images and LiDAR point clouds, as in Eq. 13.
  • The auxiliary supervisions for unlabeled data, obtained from text-driven knowledge guidance in Eq. 15.
@@ 5 Experiments
Table 2: Benchmarking results among state-of-the-art approaches using the LiDAR range view, bird’s eye view (BEV), sparse voxel, and multi-view fusion backbones. A unified backbone setup is used across representations, except for GPC [30] and LiM3D [44], which adopt extra modules. All mIoU scores are given in percentage ($\%$). The sup.-only, best, and second best scores under each data split within each group are shaded with gray, blue, and yellow, respectively.
@@
Range View
Sup.-only - FIDNet $38.3$ $57.5$ $62.7$ $67.6$ $36.2$ $52.2$ $55.9$ $57.2$ $33.1$ $47.7$ $49.9$ $52.5$
@@
Voxel
Sup.-only - Cylinder3D $50.9$ $65.9$ $66.6$ $71.2$ $45.4$ $56.1$ $57.8$ $58.7$ $39.2$ $48.0$ $52.1$ $53.8$
@@
LiM3D [44] CVPR’23 - - - - - $61.6$ $62.6$ $62.8$ - $60.3$ $60.5$ $60.9$
ImageTo360 [74] ICCVW’23 - - - - $54.1$ $60.0$ $62.2$ $65.0$ - - - -
@@
Fusion
Sup.-only - SPVCNN $57.9$ $71.7$ $73.0$ $74.6$ $52.7$ $64.1$ $64.5$ $65.1$ $47.2$ $57.3$ $58.2$ $58.8$
@@
Table 3: Robustness enhancement effect analysis among different mixing-based 3D scene augmentation methods on the Robo3D [36] benchmark for LiDAR semantic segmentation (w/ a MinkUNet [12] backbone) and 3D object detection (w/ a CenterPoint [95] backbone) tasks. All mIoU/mAP/mCE/mRR scores are given in percentage ($\%$). The baseline, best, and second best scores under each evaluation metric are shaded with gray, blue, and yellow, respectively.
@@
SemanticKITTI-C
None - MinkUNet $100.0$ $81.9$ $55.9$ $54.0$ $53.3$ $32.9$ $56.3$ $58.3$ $54.4$ $46.1$
@@ 5.2 Experimental Setups
Implementation Details. Our LaserMix++ framework is established based on the MMDetection3D codebase [13]. We denote the supervised-only baseline as sup.-only in our experiments. For semi-supervised learning, the labeled/unlabeled data are selected in four ways as shown in Figure 4: 1) random sampling, 2) uniform sampling, 3) sequential sampling, and 4) ST-RFD [44] sampling strategies. Due to the lack of previous works, we also compare consistency regularization [82, 9, 19] and entropy minimization [110] methods from semi-supervised image segmentation. The number of spatial area sets $m$ in multi-modal LaserMix is uniformly sampled from 2 to 6. The loss weights $\lambda_{\text{mix}}$, $\lambda_{\text{mt}}$, $\lambda_{\text{c2l}}$, and $\lambda_{\text{lkg}}$ in Eq. 16 are set to 2.0, 250, 1.5, and 1.0, respectively. All experiments are implemented using PyTorch on eight NVIDIA A100 GPUs. All baseline models are trained with a batch size of 2 on each GPU, along with the AdamW optimizer [56], OneCycle learning rate schedule [77], and a learning rate of 0.008. We do not include any type of test-time augmentation or model ensemble during the evaluation. For additional details, kindly refer to the Appendix.
Evaluation Metrics. In this work, we report the intersection-over-union (IoU) for each semantic class and the mean IoU (mIoU) scores across all semantic classes in semi-supervised learning benchmarks. For out-of-distribution robustness assessment, we follow the Robo3D [36] protocol to report the mean corruption error (mCE) and mean resilience rate (mRR).
5.3 Comparative Study
Improvements over Baselines. In Table 2, we compare the proposed LaserMix++ with the LaserMix [38] and sup.-only baselines on nuScenes [18], SemanticKITTI [2], and ScribbleKITTI [83]. Since our approaches can be universally applied to different LiDAR representations, we also set up the benchmark using various backbones from the literature. We observe consistent improvements achieved across all different settings by integrating multi-modal learning for data-efficient 3D scene understanding. On average, there is a 2% to 3% performance gain brought by LaserMix++ over the previous framework. The improvements are especially predominant when the number of labeled data is extremely limited (e.g., 1%), which verifies the effectiveness of our approaches. As for the sup.-only baselines, the gains vary under different backbones. LaserMix++ yields up to 13% mIoU improvements across three datasets when using the range view backbone [104]. Similar trends hold for the bird’s eye view [102], sparse voxel [107, 12], and multi-view fusion [81] backbones, where the mIoU gains under the 1% split are around 6%, 8%, and 10%, respectively. The results strongly verify the effectiveness of our framework and further highlight the importance of leveraging unlabeled data for LiDAR semantic segmentation.
Table 4: Benchmarking results among existing 3D representation learning methods pre-trained, linear-probed (LP), and fine-tuned on nuScenes [18]. All mIoU scores are given in percentage ($\%$). The sup.-only, best, and second best scores are shaded with gray, blue, and yellow, respectively.
@@ 5.3 Comparative Study
Compare with State of the Arts. We compare LaserMix++ with recent arts from the literature, i.e., GPC [30], CRB [83], and Lim3D [44], and show results (using the Cylinder3D backbone) in Table 2. As can be seen, LaserMix++ exhibits much better results than GPC [30] and [83] across various scenarios. The most recent work, Lim3D [44], achieves the best performance on ScribbleKITTI [83]. However, its results were obtained by voting from multiple augmented test-time ensembles, which created a different evaluation protocol from ours. In addition to the above, we also reproduced several popular algorithms [82, 110, 9] from the semi-supervised image segmentation domain. The results in Table 2 verify that these methods, albeit competitive in 2D tasks, only yield sub-par performance in the semi-supervised LiDAR semantic segmentation benchmark. This highlights the importance of exploiting the LiDAR data structure for data-efficient 3D scene understanding.
Table 5: Ablation study for different components in the LaserMix++ framework (w/ a FIDNet [104] backbone) on the official val sets of SemanticKITTI [2] and nuScenes [18]. Settings: (a) The sup.-only results. (b) The baseline [82] results; (c) The baseline [82] results with multi-modal operations (camera-to-LiDAR feature distillation and language-driven knowledge guidance). (d) The LaserMix results w/ Student net supervision ($S$); (e) The LaserMix results w/ Teacher net supervision (TS). (f) The LaserMix++ results w/ the camera-to-LiDAR feature distillation ($\mathcal{L}_{\text{c2l}}$). (g) The LaserMix++ results w/ the language-driven knowledge guidance ($\mathcal{L}_{\text{lkg}}$). (h) The complete LaserMix++ configuration results. All mIoU scores are given in percentage ($\%$).
# $\mathcal{L}_{\text{mt}}$ $\mathcal{L}_{\text{mix}}$ $S$ TS $\mathcal{L}_{\text{c2l}}$ $\mathcal{L}_{\text{lkg}}$ SemanticKITTI [2] nuScenes [18]
@@
[…]
Table 6: Benchmark results of different semi-supervised learning approaches on the Cityscapes [14] dataset. (a) Methods using MeanTeacher [82] as the backbone. (b) Methods using CPS [9] as the backbone with a CutMix [97] augmentation. All mIoU scores are given in percentage ($\%$).
# Method $\mathbf{1/16}$ $\mathbf{1/8}$ $\mathbf{1/4}$ $\mathbf{1/2}$
a MeanTeacher [82] $66.1$ $71.2$ $74.4$ $76.3$
CCT [67] $66.4$ $72.5$ $75.7$ $76.8$
GCT [32] $65.8$ $71.3$ $75.3$ $77.1$
CPS [9] $69.8$ $74.4$ $76.9$ $78.6$
b CPS-CutMix [9] $74.5$ $76.6$ $77.8$ $78.8$
Qualitative Assessments. Figure 5 displays visualizations of the 3D scene segmentation results for different semi-supervised learning algorithms on the validation set of nuScenes [18], where each example covers a $\times50$ $\text{m}^{2}$ driving scene centered by the ego-vehicle. We observe that previous methods can only improve predictions in limited regions, while our approaches holistically eliminate false predictions in almost every region around the ego-vehicle. The consistency enlightened by our methods has yielded better 3D segmentation accuracy under annotation scarcity.
Enhancing Representation Learning Effects. Recent explorations on self-supervised LiDAR semantic segmentation exhibit promising representation learning effects during model pretraining [86, 103, 54, 76, 58, 52]. Such methods establish suitable self-supervised pretraining objectives and probe the qualitative aspects of representation learning via few-shot fine-tuning.
The results shown in Table 4 verify that our approaches are effective in enhancing the effects of representation learning during fine-tuning. combined with Seal [52], LaserMix++ also achieves superior performance under full supervision (from 74.7% mIoU to 77.9% mIoU). This study further proves the versatility of our framework in handling different LiDAR semantic segmentation tasks.
Enhancing Out-of-Distribution Robustness. The ability to tackle scenarios that are not observed during the training time is crucial for a 3D scene understanding model, especially under the autonomous driving context [39, 87]. The fine-grained manipulations of LiDAR point clouds in LaserMix and LaserMix++ have the potential to enrich the training distribution and, in return, achieve better robustness. In this work, we conduct experiments on the Robo3D benchmark [36] to validate the robustness enhancement effect of our framework.
As shown in Table 3, our approaches consistently yield lower mCE scores for both LiDAR semantic segmentation and 3D object detection compared to other mixing-based techniques, such as Mix3D [63] and PolarMix [85].
Table 7: Ablation study on laser beam partition strategies in LaserMix. Horizontal axis: inclination direction $\phi$; Vertical axis: azimuth direction $\alpha$. A (ii-$\alpha$, $j$-$\phi$) strategy denotes that there are $i$ azimuth and $j$ inclination partitions in total.
@@
$60.4$ $63.5_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+3.1)}}}}$ $5.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+4.8)}}}} 66$$5_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+6.1)}}}} 66.$$6.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.8)}}}}$ $65.4_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.0)}}}}$
@@
$61.5_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+1.1)}}}}$ $63.3_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+2.9)}}}}$ $5.9_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.5)}}}} 66$$1_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.7)}}}} 66.$$6.7_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+6.3)}}}}$ $65.3_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+4.9)}}}}$
@@
$60.9_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+0.6)}}}}$ $64.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+3.8)}}}}$ $5.9_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.5)}}}} 66$$3_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.9)}}}} 66.$$6.0_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+5.6)}}}}$ $65.2_{\mathbf{{{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}(+4.8)}}}}$
@@
Generalize to Image Segmentation. The proposed scene prior-based consistency regularization can be extended to image domains since they also reflect real-world driving scenarios. To validate this, we conduct experiments on Cityscapes [14] and show the results in Table 6. As can be seen, our approaches achieved improved performance over the strong image segmentation baselines [82, 9, 67, 32]. Such a generalizability ensures our approaches are universally applicable to different sensor modalities.
5.4 Ablation Study
Table 8: Ablation study on state-of-the-art 3D SSL methods (w/ the same Cylinder3D [107] backbone except for LiM3D [44]) under different data splitting strategies, i.e., (1) random, (2) uniform, (3) sequential, and (4) spatio-temporal redundant frame downsampling (ST-RFD) strategies. All mIoU scores are given in percentage ($\%$). The sup.-only, best, and second best scores under each data split within each data splitting group are shaded with gray, blue, and yellow, respectively.